I am supposed to solve the system 2x-y=6 and -4x+2y=9 by addition. I am so confused. Please help me. Thank you.
Answer by stanbon(75887)

solve the system
1st: 2x-y=6 and
2nd: -4x+2y=9 by addition
-----------
Multiply 1st be 2 to get:
3rd: 4x-2y=12
Add 2nd and 3rd to get:
0=21
This is a contradiction; it means the system
is inconsistent. It means the lines are parallel
and do not share any points.
Your answer should be "No solution".
Cheers,
Stan H.